# Water-based Quantum Dots Liquid Scintillator for Particle Physics

### v2

@@ Abstract
Liquid scintillators are typically composed from organic compounds dissolved in organic solvents. However, usage of such material is often restricted due to fire safety and environmental reasons. Because of this, R&D of water-based liquid scintillators is of extreme relevance; yet, no such scintillators have been made commercially available as yet. Here, we investigate an alternative, water-based quantum dots liquid scintillator. Pre-determined and controllable optical properties of the quantum dots, as well as the existence of large libraries of established protocols for their dispersion in aqueous solutions, make them an attractive option for nuclear and particle physics applications. We characterize the optical properties of water-based quantum dots liquid scintillator and find that most of its optical properties are preserved upon quantum dots’ phase transfer into water, through the addition of an oleic acid hydrophilic layer. Using the developed scintillator, the time and charge responses from atmospheric muons are measured, highlighting the practical viability of water-based quantum dots liquid scintillators for nuclear and particle physics, with special interest in neutrino physics.
@@ 1 Introduction
Here we are investigating the possibility of utlizing QDs for particle physics applications [13, 14, 15]. In particular, we are interested in the possibility of WbQDs liquid scintillator in nuclear fission reactor neutrino detection. Water is a safe material to use near the fission reactor core, and such a detector can be a new type of near-field monitor for fission reactors [16, 17, 18, 19, 20, 21, 22]. Typical compositions of the QDs include elements such as cadmium, selenium, and tellurium, all of which have high neutron capture cross-sections. This opens a possibility to tag neutrons, and WbQDs to act as conventional liquid scintillators with neutron detection ability without doping neutron capturing materials, in contrast to the proposed gadolinium-doped water-based liquid scintillator [20].
@@
Regarding the emission properties of QDs, the vast majority of photodetectors used for radiation detection have maximum quantum efficiencies (QE) in the blue end of the spectrum; therefore, the selection of QDs’ composition and size needs to be such as to yield emission in this spectral range. QDs with Cd-based compositions and small core sizes seem appropriate for proof-of-principle demonstrations. Furthermore, and although not explored in this work, cadmium has a high neutron capture cross-section and Cd-based cores of our QDs open a possibility to employ them as neutron capture materials, allowing tagging of neutrino interactions through the inverse beta-decay processes [24]. Regardless of the application, the concentration of QDs in the scintillating liquid needs to be high enough to facilitate sufficient muon interactions or neutron capture events for experimentation, whilst low enough to maintain the optical transparency of the liquid. As such, there will be an optimum range for QD concentration in WbQD scintillating liquids.
Finally, for application purposes, QDs must remain colloidally and functionally stable for relatively long periods of time. Whilst it is known that phase-transferred QDs might gradually lose their stability over time and that routes to mitigate this do exist [25, 26, 27, 28], little is known about the effect of long-term exposure to radiation on the QDs’ stability and functionality. Although radiation responses of QDs were studied in the past [9, 10], these were for low energy particles and not MIPs such as cosmic rays. As such, measuring MIP responses and therefore demonstrating the scintillating property of WbQDs liquid scintillator - the subject the work presented here - is the first step towards the utilization of WbQDs as particle scintillators.
@@ 2.2 Surfactant encapsulation method
Three samples were prepared in this manner, one in each of the years of 2021, 2022 and 2023, and labelled accordingly. Figure 1 shows the three samples under natural and UV illumination (recorded in August of 2023). The blue color of solutions under UV illumination (in panel (b)) is indicative of QD presence in the dispersions and therefore of the successful completion of the phase transfer procedure on each occasion. It is clear that in all three cases the concentration of QDs is relatively high; yet, and as is evident from panel (a) of Figure 1, all three samples have a high degree of transparency, as per requirements for successful neutrino experiments. In the following result section, we present a representative set of data from the 2023 sample; the properties of 2021 and 2022 samples will be discussed in detail in section 3.3, where sample aging is considered.
@@ 3.1 Optical properties of CdS/ZnS QDs
The extinction and photoluminescence (PL) spectra of QD dispersions were recorded using a 2600i UV-vis spectrometer from Shimadzu [34] and Cary Eclipse Fluorescence Spectrophotometer [35] respectively. In our interpretation of the extinction data, we attributed any strong changes in the extinction of the sample to the absorption by the QDs, since they have relatively small scattering cross-sections [36]. However, below the band gap of the QDs, scattering contributions can and do become dominant.
@@
Figure 2: Optical properties of CdS/ZnS QDs. Panels (a) and (b) show absorbance (solid lines, left scales in absorbance unit or a.u.) and photoluminescence (dashed lines, right scales in arbitrary normalization) of (a) toluene- and (b) water-dispersed QDs. Panel (c) compares the normalized photoluminescence spectra of the toluene- and water-dispersed QDs samples. Panel (d) shows the experimental data (black scatter points) for the water-dispersed QD sample and the Gaussian fit to the data that was used to determine the position and width of the first excitonic peak of the QDs. The $R^{2}$ value for the fit was 0.99983 and the corresponding $\chi^{2}$ value was $1.2\times 10^{-8}$.
@@
As expected, and as shown in Figs. 2(a) and (b), the absorbance of QDs dispersions in toluene and in water remains negligible at low energies (high wavelengths), but increases rapidly as the energy of the excitation becomes greater than the bandgap of the QDs (at $\sim$470 nm for both). No significant changes to the absorption spectra could be observed for the WbQD liquid as compared to the toluene-dispersed QDs, indicating that no significant chemical or physical changes to the QD core occurred as a result of the phase transfer process. However, comparison of the extinction spectra of the two samples at wavelengths above the QDs’ band edge, it is clear that the phase transfer of QDs to an aqueous environment resulted in an increased Rayleigh’s scattering by the QDs - this appears as an elevated but gradually decreasing background at wavelengths larger than 470 nm. This is a typical observation, owing to the size dependency of the Rayleigh scattering process and simply indicates an increased overall size of the QDs upon addition of the surfactant coating that is needed to achieve water-dispersibility [37, 38].
@@
The PL spectra of both WbQD and toluene-based dispersions appeared as Gaussian peaks centred around 460 nm (Fig. 2(a) and (b)). As discussed previously, the observed emission in the blue end of the spectrum is favourable for radiation detection experimentation. The Gaussian shape of the PL spectra is reflective of the size-dispersion of the QDs in the sample, and is very typical and characteristic of colloidal QDs [39, 40]. Of note is also the lack of any significant changes to the emission of the QDs upon phaser transfer, as is evidenced by the normalized PL spectra presented in Fig. 2(c) where only a slight shift in the position of the peak could be observed ($\Delta\lambda<1~{}\mathrm{nm}$). This small shift can be attributed to the slight change in the environmental refractive index of the QDs resulting from the addition of an OA shell during the phase transfer process [41].
@@
Figure 3: Determination of PLQY of WbQD liquid. Panels (a) and (b) show photoluminescence spectra of (a) Atto390 and (b) WbQD samples of seven different concentrations. Absorbance of each sample at the excitation wavelength ($\lambda_{exc}=$380 nm) is stated in the legend of each panel. (c) Extinction spectra of Atto390 (black line) and WbQD samples (red line). The excitation wavelength used for acquisition of photoluminescence spectra, shown in panels (a) and (b), is indicated by the gray dashed line. (d) Integrated photoluminescence intensity of Atto390 (black data) and CdS QDs (red data) as a function of sample absorbance at 380 nm. The experimental data are shown as scatter points. The lines are linear fits to the data; the $R^{2}$ value of each fit and the slope obtained from it are indicated next to the lines in respective colours. The ratio of the slopes is proportional to the ratio of the quantum yields of the two samples.
@@
Clearly, there was a reduction of the PLQY of the QY upon their phase transfer from organic to aqueous environment, which is a typical observation for such processes [43, 44], often attributed to the introduction of surface defects that act as long-lived trap states for the excited states of the QDs [45]. Nonetheless, the PLQY of the WbQD solution remained sufficiently high for the proposed proof-of-principle demonstrations.
@@
In addition to information provided above, the absorption spectrum of the QD dispersions was also utilized to estimate the concentration of the prepared dispersions, using a set of empirical equations developed by W. W. Yu et al. for aqueous dispersions of CdS QDs [46]. To do so, the measured absorption spectrum of the WbQD solution was first fitted to a Gaussian function, at the on-set of QD absorption, and as shown in Figure 2(d). Specifically, a subset of the data in the $450-490$ nm spectral range was fitted to a function of the form:
@@ 3.2 Physical properties of CdS/ZnS QDs
Figure 4: Measurements of QDs’ size and zeta potential. (a) Normalized correlation curves for CdS/ZnS QDs dispersed in toluene (blue) and in water (red). (b) Intensity hydrodynamic size distribution of CdS/ZnS QDs dispersed in toluene (blue) and in water (red), obtained from DLS measurements. (c) Intensity zeta potential distributions of CdS/ZnS QDs dispersed in water, measured using DLS measurements. In panels (a)-(c), the thinner non-solid lines show the repeats of individual measurements, whereas the thicker solid lines in brighter colour show the corresponding averages. (d) TEM image of the phase-transferred water-based CdS/ZnS QDs. The scale bar is 20 nm long.
Figure 4(a) shows the normalized correlation curves of individual measurements as well as their average for CdS QDs dispersed in toluene and in water, after phase transfer. From this data, it is clear that correlation in scattered signals persists for longer lag times for WbQDs, which is indicative of their increased hydrodynamic size as compared to QDs dispersed in toluene. This was confirmed by the cumulant analysis, which yielded mean QD sizes of $25\pm 9$ nm and $71.6\pm 0.3$ nm pre- and post-phase transfer, with the population standard deviations of $11$ nm and $26$ nm respectively. The size distribution data, shown in Figure 4(b) further confirms these results - the hydrodynamic size of WbQDs increases strongly compared to their predecessors dispersed in toluene. It is important to note that the observed large increase in the size of the QDs (approx. 3 fold) is due to both their increased physical size, owing to the additional OA layer around each QD, but also due to the emergence of charges on their surfaces which interact with the surrounding polar water molecules. It is known that the electrostatic force between charged particles and the polar water molecules contributes to the friction experienced by the particle when diffusing, resulting in larger effective hydrodynamic sizes [49, 50]. We confirmed the existence of surface charges on the phase-transferred QDs by measuring their zeta potential using DLS. Figure 4(c) shows the results of these measurements for the WbQD sample. Analysis of these results (performed automatically by the Zetasizer software) yielded a mean zeta-potential value of
@@
Finally, the physical size of the QDs’ semiconductor core was determined using Transmission Electron Microscopy (TEM). A representative TEM image obtained from the measurements is shown in Fig. 4, which was analysed to obtain the mean QD diameter of $5.0\pm 1.6$ nm. This value is comparable to that obtained from Eq. 3.5, and any discrepancy between the two can be accounted for by the limited numbers of QDs that could be imaged by the TEM technique.
@@ 3.3 Effects of sample ageing
Figure 5: Effects of sample ageing on the spectral properties of phase-transferred CdSe QDs. Panels (a) and (b) show absorbance (black lines, left scales) and fluorescence (red lines, right scales) spectra of CdS QDs phase-transferred in July of (a) 2021 and (b) 2022. The top graphs in each panel show the optical spectra recorded for CdS samples immediately after the phase-transfer process, and the bottom graphs show the same set of data measured in July of 2023. The gray dashed line in each panel indicates the peak wavelength of the freshly-transferred QDs’ emission.
@@ 4.1 Setup
Figure 6: (a) Cosmic ray test diagram. 2 cosmic ray taggers are piled to define a cosmic ray trajectory. A bottle of CdS QDs in a water solution is placed in a dark box with a 3-inch PMT located in front of it. Coincidence signals from the cosmic ray taggers are used to trigger the digitizer to record PMT pulses. (b) A view of the inside of the dark box. Note, the cosmic ray tagger (the blue box) is located outside of the dark box unlike this picture.
@@
Fig. 7 shows raw data distributions. The left plot shows the timing distributions. Here, "t=0" is defined by the cosmic ray trigger signal. Since the QDs sample has only a few cm depth in a bottle, it takes $<<1$ ns for cosmic rays to penetrate it. Thus the time distribution represents decay time of QDs by excitation due to cosmic rays. The fit gives the decay time of QDs to be $\sim$6 ns, which is not atypical for a CdS QD sample  [56, 57, 58]. The right plot of Fig. 7 shows the charge distribution in the unit of $mV\times ns$. The wide charge distribution indicates a wide range in both the angles at which the cosmic rays enter the QD sample, and in their propagation distances once in the water media, resulting in deposition of a wide range of energies across the detected events. To better understand the measured data and their relation to the geometrical factors of the experimental set-up, a simulation was developed to estimate the numbers of photons of different energies that could be produced by the cosmic rays passing through our WbQD scintillating material.

### v3

@@ Abstract
Liquid scintillators are typically composed from organic compounds dissolved in organic solvents. However, usage of such material is often restricted due to fire safety and environmental reasons. Because of this, R&D of water-based liquid scintillators is of extreme relevance; yet, no such scintillators have been made commercially available as yet. Here, we investigate an alternative, water-based quantum dots liquid scintillator. Pre-determined and controllable optical properties of the quantum dots, as well as the existence of large libraries of established protocols for their dispersion in aqueous solutions, make them an attractive option for nuclear and particle physics applications. We characterize the optical properties of water-based quantum dots liquid scintillator and find that most of its optical properties are preserved upon quantum dots’ phase transfer into water, through the addition of an oleic acid hydrophilic layer. Using the developed scintillator, the time and charge responses from atmospheric muons are measured, highlighting the practical viability of water-based quantum dots liquid scintillators for nuclear and particle physics, special interest on neutrino physics.
@@ 1 Introduction
Here we are investigating the possibility of utlizing QDs for particle physics applications [13, 14, 15]. In particular, we are interested in the possibility of WbQDs liquid scintillator in nuclear fission reactor neutrino detection. Water is a safe material to use near the fission reactor core, and such detector can be a new type of near-field fission reactor monitoring detector [16, 17, 18, 19, 20]. Typical compositions of the QDs include elements such as cadmium, selenium, and tellurium; and whilst some of them pose environmental concerns [21, 22], all of them have high neutron capture cross-sections, which opens a possibility to tag neutrons. Hence, WbQDs can act as conventional liquid scintillators with neutron detection ability without requirements for doping with neutron capturing materials, in contrast to the proposed gadolinium-doped water-based liquid scintillator [20].
@@
Regarding the emission properties of QDs, the vast majority of photodetectors used for radiation detection have maximum quantum efficiencies (QE) in the blue end of the spectrum; therefore, the selection of QDs’ composition and size needs to be such as to yield emission in this spectral range. QDs with Cd-based compositions and small core sizes seem appropriate for proof-of-principle demonstrations. Furthermore, and although not explored in this work, cadmium has a high-neutron capture cross-section and Cd-based cores of our QDs open a possibility to employ them as neutron capture materials, allowing tagging of neutrino interactions through the inverse beta-decay processes [24]. Regardless of the application, the concentration of QDs in the scintillating liquid needs to be high enough to facilitate sufficient muon interactions or neutron capture events for experimentation, whilst low enough to maintain the optical transparency of the liquid. As such, there will be an optimum range for QD concentration in WbQD scintillating liquids.
Finally, for application purposes, QDs must remain colloidally and functionally stable for relatively long periods of time. Whilst it is known that phase-transferred QDs might gradually loose their stability over time and that routes to mitigate this do exist [25, 26, 27, 28], little is known about the effect of long-term exposure to radiation on the QDs’ stability and functionality. Although radiation responses of QDs were studied in the past [9, 10], these were for low energy particles and not MIPs such as cosmic rays. As such, measuring MIP responses and therefore demonstrating the scintillating property of WbQDs liquid scintillator - the subject the work presented here - is the first step towards the utilization of WbQDs as particle scintillators.
@@ 2.2 Surfactant encapsulation method
Three samples were prepared in this manner, one in each of the years of 2021, 2022 and 2023, and labelled accordingly. Figure 1 shows the three samples under natural and UV illumination (recorded in August of 2023). The blue color of solutions under UV illumination (in panel (b)) is indicative of QD presence in the dispersions and therefore of the successful completion of the phase transfer procedure on each occasion. It is clear that in all three cases the concentration of QDs is relatively high; yet, and as is evident from panel (a) of Figure 1, all three samples have a high degree of transparency, as per requirements for successful neutrino experiments. In the following result section, we present a representative set of data from the 2023 sample; the properties of 2021 and 2022 samples will be discussed in detail in the section 3.3, where sample aging is considered. We note that, post-fabrication, all samples were stored in a non-transparent case at $4^{\circ}$. Prior to testing, a small aliquot of the sample was equilibrated at room temperature. This was done to ensure that factors such as varied storage conditions and various degrees of sample agitation did not contribute to any observable experimental differences over time.
@@ 3.1 Optical properties of CdS/ZnS QDs
The extinction and photoluminescence (PL) spectra of QD dispersions were recorded using 2600i UV-vis spectrometer from Shimadzu [34] and Cary Eclipse Fluorescence Spectrophotometer [35] respectively. In our interpretation of the extinction data, we attributed any strong changes in the extinction of the sample to the absorption by the QDs, since they have relatively small scattering cross-sections [36]. However, below the band gap of the QDs, scattering contributions can and do become dominant.
@@
Figure 2: Optical properties of CdS/ZnS QDs. Panels (a) and (b) show absorbance (solid lines, left scales in absorbance unit or a.u.) and photoluminescence (dashed lines, right scales in arbitrary normalization) of (a) toluene- and (b) water-dispersed QDs. Panel (c) compares the normalized photoluminescence spectra of the toluene- and water-dispersed QDs samples, both measured at 380 nm excitation. Panel (d) shows the experimental data (black scatter points) for the water-dispersed QD sample and the Gaussian fit to the data, that was used to determine the position and width of the first excitonic peak of the QDs. The $R^{2}$ value for the fit was 0.99983 and the corresponding $\chi^{2}$ value was $1.2\times 10^{-8}$.
@@
As expected, and as shown in Figs. 2(a) and (b), the absorbance of QDs dispersions in toluene and in water remains negligible at low energies (high wavelengths), but increases rapidly as the energy of the excitation becomes greater than the bandgap of the QDs (at $\sim$470 nm for both). No significant changes to the absorption spectra could be observed for the WbQD liquid as compared to the toluene-dispersed QDs, indicating that no significant chemical or physical changes to the QD core ocurred as a result of the phase transfer process. However, comparison of the extinction spectra of the two samples at wavelengths above the QDs’ band edge, it is clear that the phase transfer of QDs to an aqueous environment resulted in an increased Rayleigh’s scattering by the QDs - this appears as an elevated but gradually decreasing background at wavelengths larger than 470 nm. This is a typical observation, owing to the size dependency of the Rayleigh scattering process and simply indicates an increased overall size of the QDs upon addition of the surfactant coating that is needed to achieve water-dispersibility [37, 38].
@@
The PL spectra of both WbQD and toluene-based dispersions appeared as Gaussian peaks centred around 460 nm (Fig. 2(a) and (b), excitation at 380 nm). As discussed previously, the observed emission in the blue end of the spectrum is favourable for radiation detection experimentation. The Gaussian shape of the PL spectra is reflective of the size-dispersion of the QDs in the sample, and is very typical and characteristic of colloidal QDs [39, 40]. Of note is also the lack of any significant changes to the emission of the QDs upon phaser transfer, as is evidenced by the normalized PL spectra presented in Fig. 2(c) where only a slight shift in the position of the peak could be observed ($\Delta\lambda<1~{}\mathrm{nm}$). This small shift can be attributed to the slight change in the environmental refractive index of the QDs resulting from the addition of an OA shell during the phase transfer process [41].
@@
Figure 3: Determination of PLQY of WbQD liquid. Panels (a) and (b) show photoluminescence spectra of (a) Atto390 and (b) WbQD samples of seven different concentrations. Absorbance of each sample at the excitation wavelength ($\lambda_{exc}=$380 nm) is stated in the legend of each panel. (c) Extinction spectra of Atto390 (black line) and WbQD samples (red line). The excitation wavelength used for acquisition of photoluminescence spectra, shown in panels (a) and (b), is indicted by the gray dashed line. (d) Integrated photoluminescence intensity of Atto390 (black data) and CdS QDs (red data) as a function of sample absorbance at 380 nm. The experimental data are shown as scatter points. The lines are linear fits to the data; the $R^{2}$ value of each fit and the slope obtained from it are indicated next to the lines in respective colours. The ratio of the slopes is proportional to the ratio of the quantum yields of the two samples.
@@
Clearly, there was an approximately 5-fold reduction of the PLQY of the QY upon their phase transfer from organic to aqueous environment, which is a typical observation for such processes [43, 44], often attributed to the introduction of surface defects that act as long-lived trap states for the excited states of the QDs [45]. Nonetheless, the PLQY of the WbQD solution remained sufficiently high for the proposed proof-of-principle demonstrations.
@@
In addition to information provided above, the absorption spectrum of the QD dispersions was also utilized to estimate the concentration of the prepared dispersions, using a set of empirical equations developed by W. W. Yu et al. for aqueous dispersions of CdS QDs [46]. It is important to note that our QDs, composed from a CdS core with ZnS shell, differ from those for which the empirical relations were derived (CdS core only). The addition of the ZnS shell to the CdS core has the effect of passivation of surface defects, leading to improved exciton confinement, quantum yields and photochemical stability of QDs [47, 48]. However, the position of the absorption and emission peaks of the QD solution remain largely unaffected by this change. Since these specific parameters weigh strongly in the method, and only an estimate for the concentration of the sample is required for discussions of results, the use of the method developed by W. W. Yu is justified on this occasion.
The implementation of the method involved the fitting of the measured absorption spectrum of the WbQD solution to a Gaussian function, at the on-set of QD absorption, and as shown in Figure 2(d). Specifically, a subset of the data in the $450-490$ nm spectral range was fitted to a function of the form:
@@ 3.2 Physical properties of CdS/ZnS QDs
Figure 4: Measurements of QDs’ size and zeta potential. (a) Normalized correlation curves for CdS/ZnS QDs dispersed in toluene (blue) and in water (red). (b) Intensity hydrodynamic size distribution of CdS/ZnS QDs dispersed in toluene (blue) and in water (red), obtained from DLS measurements. (c) Intensity zeta potential distributions of CdS/ZnS QDs dispersed in water, measured using DLS measurements. In panels (a)-(c), the thinner non-solid lines show the repeats of individual measurements, whereas the thicker solid lines in brighter colour show the corresponding averages. (d) TEM images of the water-based CdS/ZnS QDs pre-phase transfer (left half) and post-phase transfer (right half). The scale bar is 50 nm long. Panels (e) and (f) show the distribution of QD diameters, obtained from TEM data (red bars), as well the normal distribution fit to the data (solid black line). Fitting was performed using OriginPro software, which also yielded mean and standard distribution of QDs sizes for each sample.
Figure 4(a) shows the normalized correlation curves of individual measurements as well as their average for CdS QDs dispersed in toluene and in water, after phase transfer. From this data, it is clear that correlation in scattered signals persists for longer lag times for WbQDs, which is indicative of their increased hydrodynamic size as compared to QDs dispersed in toluene. This was confirmed by the cumulant analysis, which yielded mean QD sizes of $25\pm 9$ nm and $71.6\pm 0.3$ nm pre- and post-phase transfer, with the population standard deviations of $11$ nm and $26$ nm respectively. The size distribution data, shown in Figure 4(b) further confirms these results - the hydrodynamic size of WbQDs increases strongly compared to their predecessors dispersed in toluene. It is important to note that the observed large increase in the size of the QDs (approx. 3 fold) is due to both their increased physical size, owing to the additional OA layer around each QD, but also due to the emergence of charges on their surfaces which interact with the surrounding polar water molecules. It is known that the electrostatic force between charged particles and the polar water molecules contributes to the friction experienced by the particle when diffusing, resulting in larger effective hydrodynamic sizes [51, 52]. We confirmed the existence of surface charges on the phase-transfered QDs by measuring their zeta potential using DLS. Figure 4(c) shows the results of these measurements for WbQD sample. Analysis of these results (performed automatically by the Zetasizer software) yielded a mean zeta-potential value of
@@
Finally, the physical size of the QDs’ inorganic layers (core + shell) was determined using Transmission Electron Microscopy (TEM), for both pre- and post-phaser transfer samples. Representative TEM images obtained from the measurements are shown in Fig. 4(c). Analysis of TEM images yielded size distribution plots for each sample, shown in Figs. 4(e) and (f). These data were fitted to a normal distribution model (usin OriginPro in-built fitting function) to obtain the mean QD diameter of pre-phase transfer QDs to be $6.4\pm 1.0$ nm and $6.4\pm 1.2$ nm post-phase transfer. The two values were found to be equal, which is expected for the phase-transfer protocol employed  [23], which has been previously shown to cause no change in the physical size of the QD inorganic core (including any inorganic shells) but a change in the overall size of the QDs due to the increased size of the OA layer surrounding each QD. Furthermore, both values are slightly larger but are comparable to the value obtained from the empirical model above ($5.22\pm 0.01$ nm, using equation 3.8). This, again, is expected as the size of the QDs produced using the empirical model accounts only for the optically-active CdS core of the QDs and not the ZnS inorganic shell.
@@ 3.3 Effects of sample ageing
Figure 5: Effects of sample ageing on the spectral properties of phase-transferred CdSe QDs. Panels (a) and (b) show absorbance (black lines, left scales) and fluorescence (red lines, right scales) spectra of CdS QDs phase-transferred in July of (a) 2021 and (b) 2022. The top graphs in each panel show the optical spectra recorded for CdS samples immediately after phase-transfer process, and the bottom graphs show the same set of data measured in July of 2023. The gray dashed line in each panel indicates the peak wavelength of the freshly-transferred QDs’ emission.
@@ 4.1 Setup
Figure 6: (a) Cosmic ray test diagram. 2 cosmic ray taggers are piled to define a cosmic ray trajectory. A bottle of CdS QDs in water solution is placed in a dark box where 3-inch PMT is located in front of it. Coincidence signals from the cosmic ray taggers are used to trigger the digitizer to record PMT pulses. (b) A view of the inside of the dark box. Note, cosmic ray tagger (the blue box) locates outside of the dark box unlike this picture.
Figure 7: A typical PMT pulse recorded by the digitizer. Panel (a) shows the full data window and panel (b) shows the close-ups around the PMT pulse.
Fig. 7 shows a typical recorded PMT pulse. Negative PMT pulses are recorded during the data acquisition window ($\sim 4~{}\mu$s). In the offline analysis, the peak time is found by simple scanning, and charge integration is performed after the baseline subtraction.
@@
Fig. 8 shows raw data distributions. The left plot shows the timing distribution of the pulse peak with an offset. Here, "t=0" is defined by the cosmic ray trigger signal. Since the QDs sample has only a few cm depth in a bottle, it takes $<<1$ ns for cosmic rays to penetrate it. Thus the time distribution is limited by the decay time of QDs by excitation due to cosmic rays and the resolution. The Gaussian fit gives the decay time of QDs to be at most $\sim$8 ns, which is not atypical for a CdS QD sample  [58, 59, 60]. The right plot of Fig. 8 shows the charge distribution in the unit of $mV\times ns$. The wide charge distribution indicates a wide range in both the angles at which the cosmic rays enter the QD sample, and in their propagation distances once in the water media, resulting in deposition of a wide range of energies across the detected events. To better understand the measured data and their relation to the geometrical factors of the experimental set-up, a simulation was developed to estimate the numbers of photons of different energies that could be produced by the cosmic rays passing through our WbQD scintillating material.